OpenStudy (anonymous):

Instructions:Type the correct answer in the box. Use numerals instead of words. If necessary, use / for the fraction bar and ^ to indicate an exponent. The formula for the volume of a cylinder is V = πr2h. The volume of a cylinder is three times the volume of a cone with the same radius and height. If the volume of a cone with the same height as a cylinder equals the volume of the cylinder, the equation for the radius of cone R in terms of the radius of cylinder r is

OpenStudy (anonymous):

Let R and r be the radius of the cone and cylinder respectively. If the height, h, is the same value for the cone and the cylinder, then the following is true:\[\pi h r^2=\frac{1}{3} \pi h R^2 \]Solve the above for R.\[R=\sqrt{3} r \]

OpenStudy (anonymous):

thats not it read the question again

OpenStudy (anonymous):

\[\frac{1}{3} \pi h R^2\text{/.}\, R\to \sqrt{3} r \]\[\pi h r^2 \]

OpenStudy (anonymous):

cant involve pi read plz

OpenStudy (anonymous):

The radius of the cone is \[\sqrt{3} \] or 1.73205 times larger than the cylinder's radius in order that both volumes are equal to each other.